# CoD-Lite: Real-Time Diffusion-Based Generative Image Compression

### 14 Apr 2026 v1

@@ Abstract
Recent advanced diffusion methods typically derive strong generative priors by scaling diffusion transformers. However, scaling fails to generalize when adapted for real-time compression scenarios that demand lightweight models. In this paper, we explore the design of real-time and lightweight diffusion codecs by addressing two pivotal questions. First, does diffusion pre-training benefit lightweight diffusion codecs? Through systematic analysis, we find that generation-oriented pre-training is less effective at small model scales whereas compression-oriented pre-training yields consistently better performance. Second, are transformers essential? We find that while global attention is crucial for standard generation, lightweight convolutions suffice for compression-oriented diffusion when paired with distillation. Guided by these findings, we establish a one-step lightweight convolution diffusion codec that achieves real-time $60$ FPS encoding and $42$ FPS decoding at 1080p. Further enhanced by distillation and adversarial learning, the proposed codec reduces bitrate by 85% at a comparable FID to MS-ILLM, bridging the gap between generative compression and practical real-time deployment. Codes are released at https://github.com/microsoft/GenCodec/CoD_Lite.
@@ 1 Introduction
Although such large-scale models achieve impressive visual realism, they incur prohibitive decoding latency even exceeding 10 seconds. Even with advances like one-step diffusion, state-of-the-art systems such as StableCodec (Zhang et al., 2025) operate below $3$ FPS, failing to meet real-time requirements. Consequently, diffusion-based codecs remain largely impractical for latency-sensitive applications.
@@ 4.1 Visualizing the Attention Landscape
To investigate this, we visualize the attention maps of a CoD model across all $26$ layers in Figure 3. The results are striking: only $7$ layers exhibit global receptive fields, while the remaining $19$ attend exclusively to local neighborhoods, revealing a clear two-phase attention pattern.
@@ A.1 Analysis on Diffusion Pre-training at Scale
Model Architecture. We adopt the PixNerd (Wang et al., 2025a) architecture for the diffusion module. For the large-scale variant (700700M parameters), we configure the hidden dimension to $1152$, with $26$ DiT blocks and $4$ decoupled pixel head blocks. For the lightweight variant (3434M parameters), we reduce the hidden dimension to $384$, with $10$ DiT blocks and $4$ decoupled pixel head blocks. The encoder-decoder framework follows CoD (Jia et al., 2025b), where the encoder applies $16\times$ spatial downsampling. The vector quantization bottleneck employs a $4$-bit codebook ($2^{4}=16$ codes), yielding an overall bitrate of $0.0156$ bpp.
@@ A.2 Analysis on Diffusion Transformers in CoD
Ablation Study Configuration. Table 1 compares the performance of global attention, local window attention, and depth-wise convolution. For local attention, we use a window size of $3$, i.e., each token calculates attention within a $3\times 3$ window. For depth-wise convolution, we use $3\times 3$ kernels. The multi-step pre-training follows the same pipeline as our main codec (Section 6 and Appendix A.3), with the exception that PatchGAN is used instead of DMD and projected GAN during Stage II fine-tuning.
@@ A.3 Real-Time Diffusion-Based Compression
Training Schedule. We employ a progressive multi-stage training strategy on $4$ A100 GPUs. The detailed configuration for each stage is provided in Table 4. The complete pre-training pipeline requires $284$ A100 GPU hours, while fine-tuning at each target bitrate requires an additional $244$ A100 GPU hours.

### 15 Apr 2026 v2

@@ Abstract
Recent advanced diffusion methods typically derive strong generative priors by scaling diffusion transformers. However, scaling fails to generalize when adapted for real-time compression scenarios that demand lightweight models. In this paper, we explore the design of real-time and lightweight diffusion codecs by addressing two pivotal questions. First, does diffusion pre-training benefit lightweight diffusion codecs? Through systematic analysis, we find that generation-oriented pre-training is less effective at small model scales whereas compression-oriented pre-training yields consistently better performance. Second, are transformers essential? We find that while global attention is crucial for standard generation, lightweight convolutions suffice for compression-oriented diffusion when paired with distillation. Guided by these findings, we establish a one-step lightweight convolution diffusion codec that achieves real-time $60$ FPS encoding and $42$ FPS decoding at 1080p. Further enhanced by distillation and adversarial learning, the proposed codec reduces bitrate by 85% at a comparable FID to MS-ILLM, bridging the gap between generative compression and practical real-time deployment. Codes are released at https://github.com/microsoft/GenCodec/tree/main/CoD_Lite.
@@ 1 Introduction
Although such large-scale models achieve impressive visual realism, they incur prohibitive decoding latency even exceeding 10 seconds. Even with advances like one-step diffusion, state-of-the-art systems such as StableCodec (Zhang et al., 2025) operate below 33 FPS, failing to meet real-time requirements. Consequently, diffusion-based codecs remain largely impractical for latency-sensitive applications.
@@ 4.1 Visualizing the Attention Landscape
To investigate this, we visualize the attention maps of a CoD model across all $26$ layers in Figure 3. The results are striking: only 77 layers exhibit global receptive fields, while the remaining 1919 attend exclusively to local neighborhoods, revealing a clear two-phase attention pattern.
@@ A.1 Analysis on Diffusion Pre-training at Scale
Model Architecture. We adopt the PixNerd (Wang et al., 2025a) architecture for the diffusion module. For the large-scale variant (700700M parameters), we configure the hidden dimension to $1152$, with $26$ DiT blocks and 44 decoupled pixel head blocks. For the lightweight variant (3434M parameters), we reduce the hidden dimension to $384$, with $10$ DiT blocks and 44 decoupled pixel head blocks. The encoder-decoder framework follows CoD (Jia et al., 2025b), where the encoder applies $16\times$ spatial downsampling. The vector quantization bottleneck employs a 44-bit codebook ($2^{4}=16$ codes), yielding an overall bitrate of $0.0156$ bpp.
@@ A.2 Analysis on Diffusion Transformers in CoD
Ablation Study Configuration. Table 1 compares the performance of global attention, local window attention, and depth-wise convolution. For local attention, we use a window size of 33, i.e., each token calculates attention within a $3\times 3$ window. For depth-wise convolution, we use $3\times 3$ kernels. The multi-step pre-training follows the same pipeline as our main codec (Section 6 and Appendix A.3), with the exception that PatchGAN is used instead of DMD and projected GAN during Stage II fine-tuning.
@@ A.3 Real-Time Diffusion-Based Compression
Training Schedule. We employ a progressive multi-stage training strategy on 44 A100 GPUs. The detailed configuration for each stage is provided in Table 4. The complete pre-training pipeline requires $284$ A100 GPU hours, while fine-tuning at each target bitrate requires an additional $244$ A100 GPU hours.